# MalPurifier: Enhancing Android Malware Detection with Adversarial Purification against Evasion Attacks

### v1

@@ 5.5 RQ5: Enhancement Brought to Other Detectors
Experimental Setup. To evaluate the flexibility and generalization of the purification model, we also package the DAE model as a plug to protect other detectors, such as Support Vector Machine (SVM), Fully Convolutional Network (FCN), Long Short Term Memory (LSTM), and Recurrent Neural Network (RNN). In detail, We use an SVM model with the Sigmoid activation and an FCN model with 3 fully-connected hidden layers (each having 512, 256, and 128 neurons) with the ReLU activation. The LSTM model has a hidden layer with 200 neurons, a sequence length of 1, and uses the Sigmoid function as the activation. Furthermore, we build the RNN model with 3 hidden layers (each having 200 neurons) and the Sigmoid activation. All these models are tuned by the Adam optimizer with 100 epochs, batch size 128, and learning rate 0.001.
Note that the DAE-based purification model works as a plug-and-play preprocessing method, in which we do not retrain the model but directly apply it from the aforementioned experiments. In addition, we wage 6 attacks on test malware examples in an adaptive manner, including BCA, PGD-$\ell_{1}$, PGD-$\ell_{2}$, $\ell_{\infty}$, Pointwise, and StepwiseMA, with the same hyper-parameters in Section 5.4.
Results. Fig. 5 depicts the accuracy improvement of these methods equipped with the DAE-based purification model on Drebin (drawn in red) and Androzoo (drawn in blue) datasets under 6 adaptive attacks. We make three observations as follows.
First, all detectors without enhancement cannot mitigate these adaptive attacks (with accuracy $\leqslant$ 16.42% on Drebin and $\leqslant$ 8.19% on Androzoo), except for the PGD-$\ell_{2}$ attack. This is because ML-based detectors themselves are very vulnerable to these evasion attacks with small perturbations. Especially, the attack effectiveness of the Pointwise attack is 100% when adversaries wage attacks on original detectors. Nevertheless, the PGD-$\ell_{2}$ attack, if running with a large number of iterations, will produce larger perturbations that may not evade detection.
@@
Answer to RQ5: The DAE model (the core module of MalPurifier) can be flexibly transferred to other detectors without retraining. Notably, it greatly enhances the robustness against evasion attacks, even when these models are highly vulnerable to such attacks.
6 Limitations and Open Issues
@@
While MalPurifier has yielded highly promising results, it is evident that this approach possesses some limitations. We additionally conduct experiments to assess the robustness of MalPurifier against the Mimicry attack, wherein the attacker introduces perturbations to a malware sample to closely resemble a benign application. Unfortunately, MalPurifier cannot effectively resist Mimicry ($\leqslant$ 60%) and has less effectiveness as malware samples are guided by more benign samples. This is primarily because the DAE model tends to purify these samples as benign due to their similarity to benign samples, leading to their misclassification. To mitigate this issue, we believe that implementing countermeasures, such as generating a more diverse set of adversarial examples or enhancing the adversarial purification model itself, would be beneficial. We plan to explore these possibilities in our future research.
@@
While our approach may not be foolproof, we firmly believe that it substantially enhances the resistance of Android malware detection against diverse evasion attacks in a lightweight and plug-and-play manner.
@@ 7.1 ML-based Android Malware Detection
Additionally, DL-based methods [40, 47] have demonstrated remarkable capabilities. For example, Andre [48] is a hybrid representation learning approach that clusters Android malware from multiple sources and classifies them using a three-layer DNN when they behave like existing families. Qiu et. al [5] proposed a framework that extracts heterogeneous features and utilizes DNN to recognize unknown and evolving malware.
@@ 7.3 Defenses against Evasion Attacks
In addition, there are also several countermeasures to identify evasion attacks through an auxiliary model. For example, Li et. al [23] introduced a Variational AutoEncoder (VAE) to distinguish benign examples from adversarial malware according to reconstruction errors, and Li et. al [20] leverages a convex DNN model-based detector to recognize the evasion attacks. Despite not requiring retraining of the target model, the auxiliary model remains closely coupled with the malware detection model and is still unable to effectively handle sophisticated and adaptive attacks.
Different from existing studies, the proposed method involves diversified perturbations to increase its generalization capabilities against novel or unseen attacks. Furthermore, the independent training of the purification model makes it computationally efficient and easily scalable.
@@ 8 Conclusion and Future Work
We devise a plug-and-play defense method against evasion attacks for malware detection. Instead of hardening the malware detector solely, an independently training purification model is used to purify adversarial malware. In the training phase, we generate adversarial examples via injecting diversified perturbations into malware, along with protective noises added onto benign samples, to achieve a trade-off between robustness and detection accuracy. Additionally, we establish a DAE-based purification model with a customized loss function, resulting in better feature representation and robust accuracy. Experiments via two Android datasets demonstrate the soundness of the method against a set of attacks, including unseen ones.
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62202097 and 62072100, in part by the China Postdoctoral Science Foundation under Grant 2022M710677, and in part by the Jiangsu Funding Program for Excellent under Grant 2022ZB137.

### v2

@@ 5.5 RQ5: Advantage and Transferability of the Purifier
Results. Fig. 5 presents the classification accuracy of different purification methods on the Drebin and Androzoo datasets under both clean and adversarial conditions. As we can see, all methods achieve high accuracy on clean data, with minor differences among them. DAE achieves the accuracy of 97.00% and 98.57% on Drebin and Androzoo, respectively, which is slightly lower than that of GAN (97.27%) on Drebin and DM (98.89%) on Androzoo. Although DAE does not always achieve the highest accuracy on clean data, it consistently delivers stable and competitive results.
However, under adversarial attacks, DAE exhibits a clear and substantial advantage over alternative approaches. On both Drebin and Androzoo datasets, the proposed purification model consistently maintains high accuracy across all attack types, with only minimal performance degradation compared to the clean setting. In detail, On Drebin, DAE maintains accuracy $\geqslant$ 95.08% across all attack types and achieves perfect accuracy of 100% under the Salt & Pepper attack, whereas other methods such as PCA, GAN, and DM experience dramatic drops in performance, with accuracy falling to even 0% under several attacks. DAE remains highly robust on Androzoo, achieving accuracy $\geqslant$ 99.24% for all adversarial attacks, while the competing methods show significant degradation, particularly under rFGSM, $\ell_{\infty}$, and MaxMA attacks.
These results demonstrate that DAE not only provides stable and competitive performance on clean data but also offers superior robustness and generalizability against a wide range of adversarial perturbations, making it a highly effective purification model for defending Android malware detectors against diverse evasion attacks.
Moreover, Fig. 6 depicts the accuracy improvement of other classifiers equipped with the DAE model on Drebin (drawn in red) and Androzoo (drawn in blue) under 6 white-box attacks. We make three observations as follows.
First, all detectors without enhancement cannot mitigate these attacks (with accuracy $\leqslant$ 16.42% on Drebin and $\leqslant$ 8.19% on Androzoo), except for the PGD-$\ell_{2}$ attack. This is because ML-based detectors are very vulnerable to these evasion attacks with small perturbations. Especially, the attack effectiveness of the Pointwise attack is 100% when adversaries wage attacks on original detectors. Nevertheless, the PGD-$\ell_{2}$ attack, if running with a lot of iterations, will produce larger perturbations that may not evade detection.
@@
Answer to RQ5: Compared with alternative purification techniques, the DAE model achieves the best trade-off between effectiveness, robustness, and generalizability under various adversarial attacks. Furthermore, it can be flexibly transferred as a plug-and-play module to enhance the robustness of different detectors, significantly improving their resistance to evasion attacks without retraining.
5.6 RQ6: Generalizability Against Advanced Attacks
Experimental Setup. Our analysis thus far has focused on conventional perturbation-based evasion attacks. However, in real-world environments, adversaries may employ more sophisticated attacks that target structural or semantic properties [46, 47] of Android applications. To answer this research question, we evaluate MalPurifier’s performance against these advanced attacks, thus providing a more rigorous test of its generalizability.
Specifically, we utilize features based on sequences of API calls extracted from the applications’ Control Flow Graphs (CFGs), following the MaMaDroid [48] methodology. The base model for all defenses continues to be a DNN, with its architecture and hyper-parameters remaining consistent with previous experiments. For MalPurifier, we augmented its DAE training set with a small set (500 samples) of adversarial examples generated via the obfuscation-based attacks detailed in RQ2. Additionally, the step size used for generating diversified adversarial perturbations (Algorithm 1) was set to 0.0001.
To ensure a fair comparison, all attack methods draw from the same set of possible manipulations sets (e.g., component injection, permission modification). Crucially, if a perturbation leads to an application crash or packaging failure during the attack process, the attempt is considered an attack failure, ensuring that only functional adversarial examples are evaluated. We then assess the defenses against four advanced black-box attacks and one powerful white-box attack as follows:
  • Random Attack (RA): A black-box attack that iteratively and uniformly samples perturbations from the malware perturbation set to inject, and queries the target model with the perturbed sample. Our implementation follows the instructions in Ref. [49].
  • MAB [50]: A reinforcement learning-based black-box attack, originally for Windows PE malware and adapted here for Android. It formulates the attack as a multi-armed bandit (MAB) problem to balance exploitation and exploration of manipulations. We implemented it following its official code [^5] https://github.com/weisong-ucr/MAB-malware.
  • AdvDroidZero [49]: A black-box query-based evasion attack framework designed for a zero-knowledge setting. It requires no prior information about the target model’s features, architecture, or parameters. We implemented it with its official source code [^6] https://github.com/gnipping/AdvDroidZero-Access-Instructions.
  • EvadeDroid [51]: A problem-space black-box attack that iteratively injects transformations into malware. It leverages an n-gram based approach to find opcode-level similarities with benign applications, and we implemented it with its official code [^7] https://github.com/HamidBostani2021/EvadeDroid.
  • HRAT [52]: A white-box structural attack specifically targeting graph-based malware detectors. It integrates a heuristic optimization model with reinforcement learning, performing four types of graph modifications that preserve functionality in the app’s bytecode. We used its official implementation [^8] https://github.com/zacharykzhao/HRAT in our experiments.
Note that the query limit was set to 10 for query-based attacks (RA, MAB, AdvDroidZero, EvadeDroid). For HRAT, which focuses on modification counts, the maximum number of allowed modifications was 50.
TABLE IV: Accuracy of different defenses against advanced structural attacks on Drebin and Androzoo datasets.
Defense RA MAB AdvDroidZero EvadeDroid HRAT
Drebin DNN 8.075 15.71 10.40 14.38 15.36
AT-rFGSMk 15.40 13.81 11.38 19.08 16.39
AT-Adam 9.880 7.708 15.94 21.31 15.36
DNN+ 27.31 40.84 34.08 40.95 56.44
KDE 10.19 21.07 12.38 15.04 20.05
FD-VAE 64.70 70.07 60.44 73.74 61.74
PAD-SMA 47.07 56.71 50.55 60.90 70.37
MalPurifier 79.74 90.88 87.85 92.38 92.08
Androzoo DNN 7.848 15.55 16.42 15.08 10.14
AT-rFGSMk 14.08 17.08 18.08 17.07 12.38
AT-Adam 15.91 17.61 19.04 19.38 10.14
DNN+ 29.08 37.92 33.38 30.08 49.64
KDE 15.81 15.71 19.30 17.90 15.37
FD-VAE 70.09 74.41 72.82 83.51 45.52
PAD-SMA 56.38 64.05 60.60 75.37 63.55
MalPurifier 78.71 92.04 81.31 94.38 81.31
Results. Table IV presents a detailed comparison of defense mechanisms against advanced structural attacks on both Drebin and Androzoo datasets. The results reveal several important trends as follows.
First, most defenses, except MalPurifier, struggle to provide adequate protection against these sophisticated attacks. For example, adversarially trained models such as AT-rFGSMk and AT-Adam achieve only 7.71% and 13.81% accuracy under MAB attacks on Drebin dataset, respectively. These models also remains similarly low across other attacks and datasets.
Second, auxiliary defenses such as DNN+ and KDE offer only marginal improvements. For instance, DNN+ achieves up to 56.44% accuracy against HRAT on Drebin, but its performance against other attacks is much lower (e.g., 27.31% for RA and 34.08% for AdvDroidZero). KDE consistently lags behind, with accuracies $\leqslant$ 21.07% on Drebin and $\leqslant$ 19.30% on Androzoo.
Third, more advanced defenses like FD-VAE and PAD-SMA show better robustness, with FD-VAE reaching up to 73.74% and PAD-SMA up to 70.37% on Drebin dataset. However, their performance still falls short of MalPurifier, with gaps ranging from 17.76% to 37.3% across different attacks and datasets.
These results indicate that defenses such as adversarial training and auxiliary mechanisms, while effective against certain perturbation-based attacks, fail to generalize to more complex, structure-targeting threats. In contrast, MalPurifier consistently achieves the highest accuracy across all attack scenarios. For black-box query attacks like MAB and EvadeDroid, the purifier effectively denoises the manipulative queries, yielding robust accuracies $\geqslant$ 90.88% on both datasets. Even more impressively, MalPurifier demonstrates strong resilience against the white-box HRAT attack, achieving accuracies of 92.08% on Drebin and 81.31% on Androzoo, respectively.
In summary, the purification-based paradigm in MalPurifier, which projects inputs back onto the learned manifold of clean data, is highly effective at defeating both black-box and white-box attacks, including those targeting structural and semantic features. Its high performance across all attack types and datasets highlights its superior generalizability and robustness compared to existing methods.
Answer to RQ6: MalPurifier shows strong generalizability against advanced attacks targeting structural and semantic features, significantly outperforming other defenses. It maintains a robust accuracy of up to 92.38% on Drebin and 94.38% on Androzoo across all tested black-box and white-box threats.
6 Discussion and Limitation
@@
While MalPurifier demonstrates strong performance, we acknowledge several limitations and avenues for future work. We additionally conduct experiments to assess the robustness of MalPurifier against the Mimicry attack, wherein the attacker introduces perturbations to a malware sample to closely resemble a benign application. Unfortunately, MalPurifier cannot effectively resist Mimicry ($\leqslant$ 60%) and has less effectiveness as malware samples are guided by more benign samples. This is primarily because the DAE model tends to purify these samples as benign due to their similarity to benign samples, leading to their misclassification. To mitigate this issue, we believe that implementing countermeasures, such as generating a more diverse set of adversarial examples or enhancing the adversarial purification model itself, would be beneficial. We plan to explore these possibilities in our future research.
In the case of a white-box attack, where the adversary has comprehensive knowledge of both the model and the purifier, MalPurifier could potentially be circumvented. Since our approach focuses on purifying perturbations without modifying or re-training the detection model, an adaptive adversary might carefully design adversarial malware that can be free from purification. For example, the adversary might directly implant tiny malicious functions into benign software, making it challenging for the purification model to distinguish and remove them. Alternatively, an adversary could try to generate inputs that cause the DAE itself to malfunction or reconstruct the sample incorrectly in a way (e.g., output a very noisy or distorted sample) that confuses the downstream classifier. Future research could explore several avenues to bolster defenses against such adaptive threats, such as investigating multiple diverse purifiers or developing iterative purification schemes.
@@
Although the proposed approach is designed for Android malware detection, the concept to eliminate perturbations and enhance the robustness of the detection system is not limited to the Android platform. We believe that our method can be extended to a variety of different malware types (e.g., ELF or EXE binaries) by small modifications. For instance, the feature extraction process might need to be modified to suit the characteristics of the target platform. Furthermore, due to the differences in file structures, the level and distribution of adversarial perturbations will also be different. Appropriate parameter adjustments may be required for the generation of adversarial samples and the injection of protective noise in the purification mechanism in this paper. Future work could investigate the feasibility of extending this approach to different malware types and the necessary modifications to ensure its effectiveness.
While our approach may not be foolproof, we firmly believe that it substantially enhances the resistance of Android malware detection against diverse evasion attacks in a lightweight and plug-and-play manner. In addition, we believe that with further improvements and optimizations, our approach can be generalized to a wider range of scenarios, providing more effective protection against various adversarial attacks.
@@ 7.1 ML-based Android Malware Detection
Additionally, DL-based methods [40, 53] have demonstrated remarkable capabilities. For example, Andre [54] is a hybrid representation learning approach that clusters Android malware from multiple sources and classifies them using a three-layer DNN when they behave like existing families. Qiu et al. [5] proposed a framework that extracts heterogeneous features and utilizes DNN to recognize unknown and evolving malware.
@@ 7.3 Defenses against Evasion Attacks
In addition, there are also several countermeasures to identify evasion attacks through an auxiliary model. For example, Li et al. [22] introduced a Variational AutoEncoder (VAE) to distinguish benign examples from adversarial malware according to reconstruction errors, and Li et al. [19] leverages a convex DNN model-based detector to recognize the evasion attacks. Despite not requiring retraining of the target model, the auxiliary model remains closely coupled with the malware detection model and is still unable to effectively handle sophisticated and adaptive attacks.
Our work differs fundamentally from these prior efforts. In detail, MalPurifier is the first to propose a complete purification framework for Android evasion attacks that integrates a diversified perturbation strategy to handle a wide threat landscape and a protective noise mechanism to preserve benign data. This holistic approach, which directly tackles the core challenges of the malware domain, allows MalPurifier to achieve superior robustness and generalizability, as demonstrated in our extensive evaluation.
@@ 8 Conclusion and Future Work
In this paper, we addressed the critical problem of defending Android malware detectors against evasion attacks. Recognizing the unique challenges posed by the discrete feature space and diverse threat landscape, we introduced MalPurifier, a novel plug-and-play purification framework. Our key contributions include a diversified adversarial perturbation mechanism to enhance robustness against unseen attacks, and a protective noise injection strategy to maintain high accuracy on benign data. We designed a DAE-based purification model with a customized loss function that combines reconstruction and prediction objectives to optimize for both sample recovery and classification performance. Extensive experiments on two large-scale datasets against a set of evasion attacks, including perturbation-based and structure-based threats, demonstrate that MalPurifier significantly outperforms state-of-the-art defenses. The source code of MalPurifier has been publicly available at https://github.com/SEU-ProactiveSecurity-Group/MalPurifier.
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62202097 and 62072100, in part by the China Postdoctoral Science Foundation under Grant 2024T170143 and 2022M710677, and in part by the Jiangsu Funding Program for Excellent under Grant 2022ZB137.